OpenStudy (anonymous):

When solving x2 + 6x = 10 by the completing the square method, what value is added to both sides? 9 5 4 3

OpenStudy (anonymous):

You have to make it look like : \(a^2 + b^2 + 2ab\) so that it can form into \((a+b)^2\)..

OpenStudy (anonymous):

You have \(x^2\) term, that's nice.. You have also one \(6x\) term , that's more nice..

OpenStudy (anonymous):

i still don't understand

OpenStudy (anonymous):

I haven't started making you understand this yet.. :P

OpenStudy (anonymous):

See, look it here can I write your expression as: \[(x)^2 + 2 \cdot x \cdot (3) = 10\] Right?/

OpenStudy (anonymous):

i guesssss

OpenStudy (anonymous):

Please make sure your guess is a sure thing...

OpenStudy (anonymous):

yeah you can but whats the point of doing that -_-

OpenStudy (anonymous):

I am making it look like the way you can understand easily.. :P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, now can you compare it with : \[(a)^2 + 2 \cdot a \cdot b \] \[\color{green}{(x)^2 + 2 \cdot x \cdot 3 }\] Compare these two expressions and tell me what is here a and b values??

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

a = 1?

OpenStudy (anonymous):

Look for a more carefully... b=3 is superbly correct..

OpenStudy (anonymous):

a = x

OpenStudy (anonymous):

Mind blowing.. :P

OpenStudy (anonymous):

Good, a = x here and b = 3, Okay?? Now, see do you know this formula : \(a^2 + 2ab + b^2 = (a+b)^2\)

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

Wait.. Do you know how to multiply??

OpenStudy (anonymous):

yes... i've never seen that formula

OpenStudy (anonymous):

If yes then multiply them: \[(a+b) \times (a+b)\]

OpenStudy (anonymous):

Use FOIL or any other technique you know..

OpenStudy (anonymous):

(x+3)(x+3) x^2 + 3x + 3x + 9 x^2 + 6x + 9

OpenStudy (anonymous):

Oh wow, you made my work more easier.. See, in your question : you have \(x^2 + 6x\) but you have 9, right??

OpenStudy (anonymous):

*I mean you don't have 9, right??

OpenStudy (anonymous):

yeah so the answer is A?

OpenStudy (anonymous):

Wait, answer is not important, I want you to know it more..

OpenStudy (anonymous):

Now you make it complete square, you will add 9 both the sides: \(x^2 + 6x + 9 = 10 + 9\) Now itself you have calculated that \(x^2 + 6x + 9 = (x+30)^2\) So, \((x+3)^2 = 19\)

OpenStudy (anonymous):

That's not 30, but 3 there..

OpenStudy (anonymous):

ohh so its right though?

OpenStudy (anonymous):

I said, your answer is right, but do not incline yourself more towards answer, know the method or technique how to solve it, you can then get any answer of any question you face.. :)

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

And Good, you co-operated very well.. :)

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

a way that is 10x easier is to take the b value and divide it by 2 and then square it

OpenStudy (anonymous):

I want to tell that only, but you kept me away from that path.. :P

OpenStudy (anonymous):

Here I wrote : \(2 \cdot x \cdot 3\) See, here I have separated 2 from 6 which is nothing but dividing 6 by 2, and you yourself replied, what for you are writing this? :(